OpenStudy (anonymous):

-7 +/1/5a/ =2 find the equation for 2

OpenStudy (anonymous):

@rose21

OpenStudy (anonymous):

@Jalishapink

OpenStudy (anonymous):

i don't get the question x)

OpenStudy (anonymous):

-7+1/5 a =2

OpenStudy (anonymous):

ye, thats way better

OpenStudy (anonymous):

you now the answer or can u help?

OpenStudy (anonymous):

You have to isolate a/find an expression for a, right ?

OpenStudy (anonymous):

find the solution of the equation.2

OpenStudy (anonymous):

-7+1/5a=2 1/5a=9 1/a=9*5 a=1/45

OpenStudy (anonymous):

if you solve the equation